# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v2

@@ 8. Legal Document Summarisation
LDS is a specialised branch of automatic summarisation that focuses on condensing legal texts, such as court judgements, into clear and informative summaries. Unlike general text summarisation, which extracts key details without following specific formatting rules, LDS must account for the distinct structure and specialised content of legal documents. These documents often include complex details, such as article numbers, statutory language, and citations that are critical for presenting the legal arguments and decisions accurately. The natural complexity of legal texts, characterised by their extensive length and detailed internal structures such as sections, articles, and paragraphs in statutes—demands tailored summarisation techniques. This need is further emphasised by the hierarchical importance of documents based on their judicial origin, where the interpretation of texts can vary between higher and lower court opinions (Kanapala et al., 2019).
@@
LDS can be approached through extractive and abstractive methods. Extractive summarisation techniques focus on identifying and extracting the most critical sentences or phrases directly from the text, maintaining the original wording and meaning. In contrast, abstractive summarisation involves generating new sentences that paraphrase the most important information, aiming for conciseness while ensuring that the essence of the legal text is preserved.
8.1. Task
Several systems have been specifically designed to summarise legal documents. One of the first systems in this field was the Fast Legal EXpert CONsultant (FLEXICON), created by Gelbart and Smith in 1991 (Gelbart and Smith, 1991a). FLEXICON utilises a keyword-based approach (Gelbart and Smith, 1991b), scanning a database of terms to pinpoint crucial segments of text. Following this, Moens et al. (1999) introduced the SALOMON system in 1999, which employs cosine similarity to cluster similar text regions, aiming to highlight relevant topics within the documents. This method aligns with other abstraction-oriented techniques seen in the work of Erkan and Radev (2004). Another system, LetSum, devised by Farzindar and Lapalme in 2004 (Farzindar, 2004), also adopts a keyword-centric strategy but uses “cue phrases” to identify text related to specific themes such as ‘Introduction’, ‘Context’, and ‘Conclusion’. Although LetSum was fairly successful in mimicking human-generated summaries, it tended to produce summaries that were excessively lengthy.
Building on previous developments in LDS, Polsley et al. (2016) introduce Casesummarizer, a tool designed for the legal domain that pre-processes legal texts into sentences, scores them using a TF-IDF matrix from extensive legal case reports, and enhances sentence scoring by identifying entities, dates, and section headings. The tool provides a user-friendly interface with scalable summaries, lists of entities and abbreviations, and a significance heat map.
Nguyen et al. (2021) propose an RL framework to enhance deep summarisation models for the legal domain, utilising Proximal Policy Optimisation with a reward function that integrates both lexical and semantic criteria. They fine-tune an extractive summarisation backbone based on BERTSUM (Liu, 2019), employing a reward model that includes lexical, sentence, and keyword-level semantics to produce better legal summaries. Schraagen et al. (2022) apply an RL approach with a Bi-LSTM and a deep learning approach based on the BART transformer model to abstractive summarisation of the Dutch case verdict database Rechtspraak.nl, combining extractive and abstractive summarisation to retain core facts while creating concise summaries.
Zhong and Litman (2022) focus on extractive summarisation of legal case decisions, proposing an unsupervised graph-based ranking model that leverages a reweighting algorithm to utilise document structure properties. They introduce a reweighting algorithm to improve sentence selection in the HipoRank model (Dong et al., 2021). It aims to reduce redundancy and enhance the selection of argumentative sentences from underrepresented sections.
Jain et al. (2024) propose a sentence scoring approach, DCESumm, which combines supervised sentence-level summary relevance prediction with unsupervised clustering-based document-level score enhancement. It utilises a Legal BERT-based Multi-Layer Perceptron model to predict the summary relevance of each sentence, refining scores through deep embedded sentence clustering to enhance the selection process by considering the global context of the document.
Liu et al. (2024) present Common Law Court Judgement Summarisation (CLSum), a pioneering dataset for summarising multi-jurisdictional common law court judgements, leveraging LLMs for data augmentation, summary generation, and evaluation. They employ a two-stage summarisation process with techniques, such as sparse attention mechanisms and efficient training methods to process lengthy legal documents within limited computational resources.
@@ 8.2. Datasets
Shen et al. (2022) introduce Multi-LexSum, an abstractive summarisation dataset tailored for U.S. federal civil rights lawsuits, containing 40,000 source documents and 9,000 expert-written summaries of diverse lengths, providing a rich resource for testing summarisation models.
Liu et al. (2024) publish CLSum, a dataset designed for summarising multi-jurisdictional common law court judgements from Australia, Hong Kong, the United Kingdom, and Canada. This dataset leverages LLMs for data augmentation and incorporates legal knowledge to enhance summary generation and evaluation. This dataset addresses the challenge of sparse labelled data in legal domains. CLSum includes a collection of judgements and summaries from prominent court websites. It employs novel techniques to enrich training sets and improve model performance in few-shot and zero-shot learning scenarios.
@@ 9. Legal Named Entity Recognition
NER is a fundamental task in NLP that involves identifying specific segments of text and classifying them into predefined categories such as ‘organisation’, ‘person’, and ‘location’ (Layegh et al., 2023). In the legal domain, NER extends to specialised recognition tasks that focus on extracting entities unique to legal texts, such as laws, legal norms, and procedural terms. This specialised form of NER is crucial for structuring legal documents and enhancing legal IR systems. Unlike general NER systems that handle common entity types, legal NER must navigate the complex language and structured format of legal documents, underscoring the need for systems and methodologies specifically tailored to the legal context.
9.1. Task
@@
Dozier et al. (2010) present pioneering work in NER within legal documents such as US case law and pleadings, employing three methodologies: lookup, contextual rules, and statistical models to detect entities, such as judges, attorneys, and legal terms. Their system adapts these approaches to the specialised context of legal texts, processing various types of documents and extracting legal entities. This work highlights the challenges and necessary adaptations for deploying NER in the legal domain, where the specialised language and high accuracy are required for successful legal analysis.
@@
Păis et al. (2021) develop a NER model for the legal domain that integrates Bi-LSTM cells and a Conditional Random Fields (CRFs) layer, utilising multiple data sources and embedding types. Their system architecture combines word embeddings from pre-trained models, character embeddings, gazetteer resources from the GeoNames [^7] https://www.geonames.org/ database and JRC-Names (Steinberger et al., 2011), and known affixes to enrich the model’s understanding of legal text. Their training process involved fine-tuning the word embeddings, while dynamically learning character embeddings with subsequent Bi-LSTM layers, enhancing the model’s capability to generalise across unseen texts. They implemented the system using a modified version of NeuroNER (Dernoncourt et al., 2017), allowing for online model serving and incorporating features, such as dropout for regularisation and gradient clipping to handle exploding gradients. They also explored ensemble methods to improve model accuracy by combining results from multiple model configurations, measuring the effectiveness through precision, recall, and F1 scores against a gold standard corpus.
@@
Smădu et al. (2022) explore domain adaptation in legal NER, focusing on the Romanian and German languages. They utilise a model architecture that integrates a pre-trained BERT layer for feature extraction with Bi-LSTM networks to handle sequence dependencies and CRFs for sequence tagging. Their approach employs domain adaptation techniques through a gradient reversal layer connected to a domain discriminator, aimed at reducing domain-specific biases and enhancing feature transferability across domains. This model trains on both legal and general domains, adapting to the peculiarities of each through adversarial learning that modifies the learning process dynamically. This setup allows for improved generalisation of the NER system across varied linguistic and domain contexts, though the actual performance varied with minimal improvements noted for German and decreased performance for the Romanian legal dataset.
@@ 9.2. Datasets
Leitner et al. (2020) present a dataset for NER focused on German federal court decisions, containing approximately 67,000 sentences and over two million tokens. This dataset features 54,000 manually annotated entities distributed across 19 fine-grained semantic classes, specifically tailored to the legal domain, such as court, judge, lawyer, law, person, and legal literature, along with over 35,000 TimeML-based time expressions. The annotations cover both broad categories, such as location, person, and organisation, and more specialised ones, such as legal norms and case-by-case regulations, distinguishing between different types of legal acts and literature. This dataset’s annotation process involved multiple cycles to refine the tagging guidelines and enhance annotation quality.
Păis et al. (2021) introduce the LegalNERo corpus, a manually annotated resource for NER in the Romanian legal domain, featuring 370 legal documents annotated with five general entity types: person, location, organisation, time expressions, and legal references. This corpus was developed to support both specific legal domain NER tasks and more general NER applications by enabling compatibility with existing general-purpose NER systems. The corpus includes rich entity annotations, with legal references showing the highest token count per entity, indicating their complexity and length. The detailed annotation process, including inter-annotator agreement assessed by Cohen’s Kappa, and the subsequent mapping of entities to RDF format, highlights the corpus’s utility and precision for advancing NER research and applications within the legal domain.
Au et al. (2022) introduce the E-NER dataset, an annotated collection derived from the US Securities and Exchange Commission’s EDGAR filings, designed for legal NER. This dataset contains filings that are rich in text, such as quarterly reports and significant event announcements, from which sentences were extracted and annotated with seven named entity classes more tailored to legal content than those in the standard CoNLL dataset (Tjong Kim Sang and De Meulder, 2003). The entities include person, location, organisation, government, court, business, and legislation/act, adjusting the CoNLL classes to better suit legal documents. E-NER contains longer sentences compared to CoNLL and includes detailed annotations of financial entities from legal company filings.
Kalamkar et al. (2022) present a corpus aimed at enhancing legal NER, containing 46,545 entities across 14 types identified in Indian High Court and Supreme Court Judgements. This corpus, split into preamble and judgement sections, includes diverse entity types detailed in their legal NER taxonomy, such as court, petitioner, respondent, and statute, among others. The training set, drawn from judgements between 1950 and 2017, features 29,964 entities and the development and test sets, spanning 2018 to 2022. This dataset not only facilitates training and evaluation of NER models specific to the legal domain but also provides a structured framework for assessing the performance of NER systems on legal texts. Their approach leverages a combination of manual annotation and ML techniques to ensure the precision of entity recognition in legal judgements.
10. Legal Argument Mining
Legal argument mining is the use of natural language processing to identify and extract arguments from legal documents. It aims to automate the extraction of argumentative content, thereby improving performance in legal research and practice. Current research focuses on developing domain-specific annotation schemes and employing advanced deep learning models to tackle the complexities of legal language and argument structures.
10.1. Task
Palau and Moens (2009) present pioneering research in LAM, focusing on detecting, classifying, and structuring arguments within legal texts. Their work introduces methods to automatically identify arguments, distinguish argumentative from non-argumentative sentences, and classify argumentative propositions as either premises or conclusions using statistical classifiers such as maximum entropy models, Naive Bayes classifiers, and SVM. The authors utilise the Araucaria and the ECHR corpora for evaluation. Results demonstrate that statistical methods can distinguish argumentative sentences, achieving approximately 80% accuracy on the ECHR corpus. Additionally, they discuss methods to resolve argument segmentation challenges through structural and semantic analyses, proposing that semantic relatedness measures (based on ontology or corpus-derived statistics) can enhance argument boundary detection. Finally, they investigate detecting argument structures through rhetorical pattern analysis and suggest employing context-free grammars as an initial step toward full argumentative parsing.
Zhang et al. (2023b) propose a graph-based framework for LAM that replaces the traditional pipeline approach with an end-to-end architecture. By modelling each legal document as a graph—where nodes represent text segments and edges capture sequential or semantic relationships—they mitigate error propagation across sub-tasks. Their method employs virtual node graph augmentation, which adds a global node connected to all text segments, and a collective classification algorithm that iteratively refines predictions using neighbouring node labels. They evaluate on the ECHR and the Court of Justice of the European Union (CJEU) datasets using Graph Convolutional Network (GCN) and Residual Gated Graph ConvNet (ResGCN) models. In particular, ResGCN demonstrates better performance on both datasets, surpassing baseline methods in classifying premises, conclusions, and non-argumentative text.
Santin et al. (2023) introduce a novel approach for predicting argumentation structures in CJEU decisions on fiscal state aid. They propose a tailored model, ResAttArg, which integrates residual connections, attention mechanisms, and multi-task learning to jointly predict argument component types and their inferential links. The model leverages domain-specific features—such as pre-trained word embeddings and encoding of inter-component distances—to capture complex argumentative relations. In contrast to a fine-tuned transformer baseline, the ensemble of ResAttArg networks shows improvements in linking argumentative components. The authors also observe that conventional oversampling strategies can degrade performance and that expanding the candidate span does not yield additional benefits. In addition, this study expand their previous dataset, Demosthenes, which consists of 40 decisions from CJEU on fiscal state aid. The corpus is annotated with detailed argumentative components—such as premises and conclusions—and is enriched with an annotation layer that captures inferential links and relation subtypes.
Habernal et al. (2024) utilise advanced NLP techniques to mine legal arguments from ECHR decisions. Specifically, they employ a multi-task transformer-based model, which is a deep learning approach designed for both argument identification and classification. This model leverages the capabilities of transformers—a widely used architecture in modern NLP—to process and understand complex legal texts, outperforming previous legal NLP models according to expert evaluations.
10.2. Datasets
@@
Poudyal et al. (2020) present the ECHR corpus for LAM, a structured dataset of 42 decisions from ECHR, annotated with argumentative components: premises, conclusions, and non-argumentative text. The corpus facilitates research on argument mining by enabling three key tasks: argument clause recognition, clause relation prediction, and premise/conclusion classification. The annotation process involved legal experts, with iterative refinements leading to 80% inter-annotator agreement. Habernal et al. (2024) work on the same dataset and expand it to 373 decisions related to Articles 3, 7, and 8 of the ECHR, totalling 2.3 million tokens with 15,205 annotated argument spans. Annotated by six law students on the INCepTION [^8] https://inception-project.github.io platform under expert supervision, it achieved high reliability (Krippendorff’s alpha $\approx$ 0.80).
11. Large Legal Corpora

### v3

@@ 8. Legal Document Summarisation
LDS is a specialised branch of automatic summarisation that condenses legal texts, such as court judgements, into clear and informative summaries. Unlike general text summarisation, which extracts key details without following specific formatting rules, LDS must accommodate the distinct structure and specialised content of legal documents. These documents often include complex elements that are essential for presenting the legal arguments and decisions accurately, such as article numbers, statutory language, and citations. The inherent complexity of legal texts, characterised by their length and detailed internal structures (sections, articles, and paragraphs), demands customised summarisation techniques. This requirement is reinforced by the hierarchical significance of documents based on judicial origin, where interpretations may differ between higher and lower court opinions (Kanapala et al., 2019).
@@
LDS can be approached through extractive and abstractive methods. Extractive techniques identify and select the most important sentences or phrases directly from the text, preserving original wording and meaning. Abstractive methods, by contrast, generate new sentences that paraphrase the key information, aiming for conciseness while maintaining the essence of the legal text.
@@ 8.1. Datasets
Shen et al. (2022) introduce Multi-LexSum, an abstractive summarisation dataset tailored for the US federal civil rights lawsuits, containing 40,000 source documents and 9,000 expert-written summaries of diverse lengths.
Liu et al. (2024) published the Common Law Court Judgement Summarisation (CLSum), a dataset designed for summarising multi-jurisdictional common law court judgements from Australia, Hong Kong, the United Kingdom, and Canada. This dataset utilises LLMs for data augmentation and incorporates legal knowledge to enhance summary generation and evaluation. This dataset addresses the challenge of sparse labelled data in legal domains. CLSum includes a collection of judgements and summaries from prominent court websites. They employ a two-stage summarisation process with techniques such as sparse attention mechanisms and efficient training methods to process lengthy legal documents with limited computational resources.
8.2. Approaches
Several systems have been specifically designed to summarise legal documents. One of the first systems in this field was the Fast Legal EXpert CONsultant (FLEXICON), created by Gelbart and Smith in 1991 (Gelbart and Smith, 1991a). FLEXICON utilises a keyword-based approach (Gelbart and Smith, 1991b), scanning a database of terms to pinpoint crucial segments of text. Following this, Moens et al. (1999) introduced the SALOMON system in 1999, which employs cosine similarity to cluster similar text regions, aiming to highlight relevant topics within the documents. This method aligns with other abstraction-oriented techniques seen in the work of Erkan and Radev (2004). LetSum (Farzindar and Lapalme, 2004), developed by Farzindar and Lapalme in 2004, also adopts a keyword-centric strategy but uses “cue phrases” to identify text related to specific themes such as ‘Introduction’, ‘Context’ and ‘Conclusion’. Although LetSum approximated human-written summaries reasonably well, it often produced documents that were longer than desired.
Building on previous developments in LDS, Polsley et al. (2016) introduce Casesummarizer, a tool designed for the legal domain that pre-processes legal texts into sentences, scores them using a TF-IDF matrix from extensive legal case reports and enhances sentence scoring by identifying entities, dates and section headings. The tool provides a user-friendly interface with scalable summaries, lists of entities and abbreviations and a significance heat map.
Nguyen et al. (2021) propose an RL framework to enhance deep summarisation models for the legal domain, utilising Proximal Policy Optimisation with a reward function that integrates both lexical and semantic criteria. They fine-tune an extractive summarisation backbone based on BERTSUM (Liu, 2019), employing a reward model that includes lexical, sentence, and keyword-level semantics to produce better legal summaries. Schraagen et al. (2022) apply an RL approach with a Bi-LSTM and a DL approach based on the BART transformer model to abstractive summarisation of the Dutch case verdict database Rechtspraak.nl, combining extractive and abstractive summarisation to retain core facts while creating concise summaries.
Zhong and Litman (2022) focus on extractive summarisation of legal case decisions, proposing an unsupervised graph-based ranking model that leverages a re-weighting algorithm to utilise document structure properties. They extend the HipoRank model (Dong et al., 2021) with a novel re-weighting algorithm to improve sentence selection, reducing redundancy and enhancing the inclusion of argumentative sentences from underrepresented sections.
Jain et al. (2024) propose a sentence scoring approach, DCESumm, which combines supervised sentence-level summary relevance prediction with unsupervised clustering-based document-level score enhancement. It utilises a Legal BERT-based Multi-Layer Perceptron model to estimate the summary relevance of each sentence, then refines these scores through deep embedded sentence clustering to incorporate the document’s global context.
@@ 9. Legal Named Entity Recognition
NER identifies and categorises textual mentions into predefined types such as organisations, persons and locations (Layegh et al., 2023). In the legal domain, NER extends to specialised recognition tasks that focus on extracting entities unique to legal texts, such as laws, legal norms, and procedural terms. This specialised form of NER is crucial for structuring legal documents and enhancing legal IR systems. Unlike general NER systems that handle common entity types, legal NER must navigate the complex language and structured format of legal documents, motivating the need for systems and methodologies specifically tailored to the legal context.
9.1. Datasets
Leitner et al. (2020) release a German legal NER corpus drawn from federal court decisions, comprising about 67,000 sentences and more than two million tokens. It contains roughly 54,000 manually annotated entities across 19 fine-grained, domain-specific classes, such as court, judge, lawyer, law, person and legal literature, alongside over 35,000 TimeML-based (Pustejovsky et al., 2003) time expressions. The annotations cover both broad categories, such as location, person and organisation and more specialised ones, such as legal norms and case-by-case regulations, distinguishing between different types of legal acts and literature. Annotation proceeded through multiple iterations to refine guidelines and ensure high-quality labels.
Păis et al. (2021) introduce the LegalNERo corpus, a manually annotated resource for NER in the Romanian legal domain, featuring 370 legal documents annotated with five entity types: person, location, organisation, time expressions and legal references. This corpus was developed to support both specific legal domain NER tasks and more general NER applications by enabling compatibility with existing general-purpose NER systems. The corpus includes rich entity annotations, with legal references showing the highest token count per entity, indicating their complexity and length. The annotation workflow involved several refinement cycles, inter-annotator agreement measured by Cohen’s kappa and conversion of entities to RDF, ensuring accuracy and usefulness for legal NER research.
Au et al. (2022) present the E-NER dataset, an annotated collection derived from the US Securities and Exchange Commission’s EDGAR filings, designed for legal NER. This dataset contains filings that are rich in text, such as quarterly reports and significant event announcements, from which sentences were extracted and annotated with seven named entity classes more tailored to legal content than those in the standard CoNLL dataset (Tjong Kim Sang and De Meulder, 2003). The entities include person, location, organisation, government, court, business and legislation/act, adjusting the CoNLL classes to better suit legal documents. E-NER contains longer sentences compared to CoNLL and includes detailed annotations of financial entities from legal company filings.
Kalamkar et al. (2022) release a legal NER corpus with 46,545 entities of 14 types extracted from Indian High Court and Supreme Court judgements. The corpus is divided into preamble and judgement sections and covers entities such as court, petitioner, respondent, and statute. The training set, drawn from judgements between 1950 and 2017, contains 29,964 entities, while the development and test sets cover cases from 2018 to 2022, ensuring no training leakage. This dataset not only facilitates training and evaluation of NER models specific to the legal domain but also provides a structured framework for assessing the performance of NER systems on legal texts. Their approach leverages a combination of manual annotation and ML techniques to ensure the precision of entity recognition in legal judgements.
9.2. Approaches
@@
Dozier et al. (2010) conduct early research on NER in legal texts, including US case law and pleadings, by combining lookup methods, contextual rules and statistical models to identify entities such as judges, attorneys and legal terms. Their system adapts these approaches to the specialised context of legal texts, processing various types of documents and extracting legal entities. This work highlights the challenges and necessary adaptations for deploying NER in the legal domain, where the specialised language and high accuracy are required for successful legal analysis.
@@
Păis et al. (2021) develop a legal NER model that combines Bi-LSTM layers with a Conditional Random Field (CRF) output layer and leverages multiple data sources and embedding types. The architecture integrates pre-trained word embeddings, character embeddings, and gazetteer entries from GeoNames [^7] https://www.geonames.org/ and JRC-Names (Steinberger et al., 2011), along with known legal affixes, to enrich text representations. During training, word embeddings are fine-tuned while character embeddings are learned dynamically through the Bi-LSTM layers, improving generalisation to unseen texts. Built on a modified version of NeuroNER (Dernoncourt et al., 2017), the system supports online serving and employs dropout for regularisation and gradient clipping to mitigate exploding gradients. The authors also explore ensembles of different model configurations, evaluating performance via precision, recall and F1 scores against a gold-standard corpus.
@@
Smădu et al. (2022) explore domain adaptation in legal NER, focusing on the Romanian and German languages. Their architecture combines a pre-trained BERT layer for feature extraction with Bi-LSTM networks to handle sequence dependencies and CRFs for sequence tagging. Their approach employs domain adaptation techniques through a gradient reversal layer connected to a domain discriminator, aimed at reducing domain-specific biases and enhancing feature transferability across domains. The model is trained on both legal and general corpora through adversarial learning, aiming to improve transferability across domains. Results show marginal gains for German but a performance drop on the Romanian legal dataset, indicating that benefits vary by language and domain.
@@
10. Legal Argument Mining
LAM applies NLP to identify and extract arguments from legal documents, automating the detection of claims, premises, and their interrelations to enhance legal research and practice. By reconstructing both the local structure of individual arguments and the global network of relations between them, it supports legal reasoning, exposing chains of reasoning that inform judicial decisions and support tasks such as conflict resolution and why-question answering (Palau and Moens, 2009). To meet these demands, recent work focuses on domain-specific annotation schemes and advanced DL models capable of handling the intricacies of legal language and argument structure. Effective argument mining therefore relies on identifying elementary argumentative units, modelling their rhetorical relations and determining whether these structures can be derived automatically.
@@ 10.1. Datasets
Poudyal et al. (2020) present the ECHR corpus for LAM, a structured dataset of 42 decisions from ECHR, annotated with argumentative components: premises, conclusions and non-argumentative text. The corpus facilitates research on argument mining by enabling three key tasks: argument clause recognition, clause relation prediction and premise/conclusion classification. The annotation process involved legal experts, with iterative refinements leading to 80% inter-annotator agreement. Habernal et al. (2024) develop an annotation scheme for ECHR judgements, grounded in legal argumentation theory and designed to capture argument spans such as claims, premises and their interrelations. Using this scheme, the authors compile and manually annotate a large corpus of 373 ECHR decisions, comprising approximately 2.3 million tokens and over 15,000 argument spans. Six law students performed the annotation under expert supervision and achieved high inter-rater agreement (as measured via Krippendorff’s alpha, with values close to 0.80).
Grundler et al. (2022) introduce Demosthenes, a corpus of 40 Court of Justice of the European Union (CJEU) decisions on fiscal state aid, annotated at the sentence level for argument mining. Each decision (written in English and spanning 2000–2018) was obtained from EUR-Lex and manually annotated to capture its argumentative reasoning, focusing on the “Findings of the Court” section where the judgement’s legal arguments are laid out. Using an iteratively refined annotation guideline, two experts with legal domain expertise labelled each sentence in this section as an argumentative element (premise or conclusion), further denoting each premise as legal or factual and assigned each argument to a category in a legal argumentation scheme typology.
10.2. Approaches
Palau and Moens (2009) present pioneering research in LAM, focusing on detecting, classifying and structuring arguments within legal texts. Their work introduces methods to automatically identify arguments, distinguish argumentative from non-argumentative sentences and classify argumentative propositions as either premises or conclusions using statistical classifiers such as maximum entropy models, Naive Bayes classifiers and SVM. The authors utilise the Araucaria and the ECHR corpora for evaluation. Results demonstrate that statistical methods can distinguish argumentative sentences, achieving approximately 80% accuracy on the ECHR corpus. Additionally, they discuss methods to resolve argument segmentation challenges through structural and semantic analyses, proposing that semantic relatedness measures (based on ontology or corpus-derived statistics) can enhance argument boundary detection. Finally, they investigate detecting argument structures through rhetorical pattern analysis and suggest employing context-free grammars as an initial step toward full argumentative parsing.
Zhang et al. (2023b) propose a graph-based framework for LAM that replaces the traditional pipeline approach with an end-to-end architecture. By modelling each legal document as a graph – where nodes represent text segments and edges capture sequential or semantic relationships – they mitigate error propagation across sub-tasks. Their method employs virtual node graph augmentation, which adds a global node connected to all text segments and a collective classification algorithm that iteratively refines predictions using neighbouring node labels. They evaluate on the ECHR and the CJEU datasets using Graph Convolutional Network (GCN) and Residual Gated GCN (ResGCN) models. In particular, ResGCN demonstrates better performance on both datasets, surpassing baseline methods in classifying premises, conclusions and non-argumentative text.
Santin et al. (2023) propose a novel annotation scheme for predicting argument structures in CJEU fiscal state aid decisions, addressing the scarcity of annotated resources and the complexity of legal reasoning. Building on the Demosthenes corpus, they refine their previous dataset by distinguishing a richer set of inferential relations, including direct support, indirect support (support from failure), rebuttal, undercut and rephrase links, which captures the logical and discursive connections support judicial arguments. Using the extended dataset, they conducted an empirical study comparing DistilRoBERTa (Sanh et al., 2019) with an ensemble of attentive residual networks  (Galassi et al., 2023) (ResAttArg) for link prediction, exploring variations in training (with/without oversampling and different link distance thresholds). The ResAttArg ensemble outperforms the distilled transformer and does so with lower computational demand.
Habernal et al. (2024) utilise advanced NLP techniques to mine legal arguments from ECHR decisions. Specifically, they employ a multi-task transformer-based model, which is a DL approach designed for both argument identification and classification. This model leverages the capabilities of transformers to process and understand complex legal texts, outperforming previous legal NLP models according to expert evaluations.
Table 4 maps each dataset to every cited study that employs it, covering both the current task and other legal NLP tasks discussed in the survey.
11. Large Legal Datasets